Identify the conic section whose equation is

5x^2-4xy+5y^2=4
1
Expert's answer
2020-02-17T12:25:17-0500

Answer: Ellipse


"5x^2 - 4xy + 5y^2 = 4"

If we draw a graph of this function we'll get:



Proof:

"\begin{Bmatrix}
 x=x_1cos(\phi)-y_1sin(\phi) \\
 y=x_1sin(\phi)+y_1cos(\phi)
\end{Bmatrix}"


"5(x_1cos(\phi)-y_1sin(\phi))^2-4(x_1cos(\phi)-y_1sin(\phi))(x_1sin(\phi)+y_1cos(\phi))+5(x_1sin(\phi)+y_1cos(\phi))^2=4"


After simplifying we will write only expression with "x_1y_1" and this expression equal to zero:

"4(x_1y_1cos^2(\phi) - x_1y_1sin^2(\phi))=0"

"\cos^2(\phi)-\sin^2(\phi)=0;\\\phi=\pi/4;"


"\begin{Bmatrix}
 x=x_1cos(\pi/4)-y_1sin(\pi/4) =\frac{x_1}{\sqrt{2}} - \frac{y_1}{\sqrt{2}} \\
 y=x_1sin(\pi/4)+y_1cos(\pi/4)=\frac{x_1}{\sqrt{2}} + \frac{y_1}{\sqrt{2}}
\end{Bmatrix}"


and we will get:


"\frac{3x_1^2}{4} + \frac{7y_1^2}{4} =1"

This is an equation of the ellipse.